Find the value of a and b so that (x+1) and (x-1)are the factors ofx4+ax3-3x2+2x+b.
1
Expert's answer
2012-06-12T07:37:35-0400
if x+1 and x-1 are factors of f(x) then f(1)=f(-1)=0

So after
substitution -1 and 1 we'll
get:
1-a-3-2+b=0
1+a-3+2+b=0

Then
b-a=4
b+a=0

Finally,
b=2, a=-2